### v2

@@ 3.2 Matrix square root
$+\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|h_{-x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
$\leq$ $\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\epsilon}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
@@
$+\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|.$
Since the integral on the small circular arc tends to 00 as $\epsilon\to 0$, we only need to take care of the integral along $\mathcal{C}^{+}_{\delta}$, $\mathcal{C}^{-}_{\delta}$ and over the large circular arc $\mathcal{C}_{R}$. On the large circular arc $\mathcal{C}_{R}$, we start with the expression:
$\left\|(M-sI)^{-1}\right\|$ $=\frac{1}{\sqrt{\lambda_{\min}\left((M-sI)^{*}(M-sI)\right)}}$
$=\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M-s^{*}M-sM^{*}+|s|^{2}I\right)}}$
(12) $\leq\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M\right)+\lambda_{\min}\left(|s|^{2}I\right)+\lambda_{\min}\left(-s^{*}M-sM^{*}\right)}}.$
Letting $T=-s^{*}M-sM^{*}$, we get
$\bm{x}^{*}T\bm{x}$ $=-s^{*}\bm{x}^{*}M\bm{x}-s\bm{x}^{*}M^{*}\bm{x}$
$=-s^{*}\bm{x}^{*}M\bm{x}-s(\bm{x}^{*}M\bm{x})^{*}.$
Thus, one obtains $\lambda_{\min}(T)=\min\limits_{\|\bm{x}\|=1}\bm{x}^{*}T\bm{x}=-2\max\limits_{\|\bm{x}\|=1}\operatorname{Re}(s^{*}\bm{x}^{*}M\bm{x}).$ In addition, applying the Cauchy-Schwarz inquality leads to
$\operatorname{Re}(s^{*}\bm{x}^{*}M\bm{x})\leq|s^{*}\bm{x}^{*}M\bm{x}|=|s|\cdot|\bm{x}^{*}M\bm{x}|\leq|s|\cdot\|M\|=|s|\sigma_{\max}(M).$
Then, we find $\lambda_{\min}(T)\geq-2|s|\sigma_{\max}(M)$. Substituting the inequality into (3.9), we have
$\left\|(M-sI)^{-1}\right\|$ $\leq\frac{1}{\sqrt{(\sigma_{\min}(M))^{2}+|s|^{2}-2|s|\sigma_{\max}(M)}}$
$=\frac{1}{\sigma_{\min}(M)}.$
Thus, over the the large circular arc $\mathcal{C}_{R}$, we obtain the following estimate:
$\frac{1}{2\pi}\int_{\mathcal{C}_{R}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\||\mathrm{d}s|\leq$ $\frac{2\pi R}{2\pi\sigma_{\min}(M)}\frac{R}{\sqrt{R^{2}+x^{2}}}$
$=$ $\frac{4\sigma_{\max}(M)\kappa(M)}{(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}.$
Along the lines $\mathcal{C}^{+}_{\delta}$, $\mathcal{C}^{-}_{\delta}$, one arrives at
$\left\|(M-sI)^{-1}\right\|$ $=\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M-s^{*}M-sM^{*}+|s|^{2}I\right)}}$
(13) $\leq\frac{1}{\sqrt{\lambda_{\min}\left(M^{*}M\right)+\lambda_{\min}(-s^{*}M-sM^{*})}}.$
Given that $M$ is positive-definite, $\frac{M+M^{*}}{2}$ is also positive-definite. We assume $s$ takes the form: $s=s_{1}\pm i\delta$. Since $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$ with $C_{M}=\frac{\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)}{\left|\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)\right|}$, then $-R\leq s_{1}\leq-\frac{2\epsilon}{\sqrt{C_{M}^{2}+4}}<-\epsilon$. Therefore,
$\lambda_{\min}(-s^{*}M-sM^{*})$ $=-2\lambda_{\max}\left(s_{1}\left(\frac{M+M^{*}}{2}\right)-i\delta\left(\frac{M-M^{*}}{2}\right)\right)$
$\geq-2\lambda_{\max}\left(s_{1}\left(\frac{M+M^{*}}{2}\right)\right)-2\lambda_{\max}\left(-i\delta\left(\frac{M-M^{*}}{2}\right)\right)$
$=-2s_{1}\lambda_{\min}\left(\frac{M+M^{*}}{2}\right)+2\delta\lambda_{\min}\left(i\left(\frac{M-M^{*}}{2}\right)\right)$
$\geq 0.$
Substituting this into 3.9 yields
$\left\|(M-sI)^{-1}\right\|\leq\frac{1}{\sigma_{\min}(M)}.$
Analogously, with $\delta=\frac{\epsilon C_{M}}{\sqrt{C_{M}^{2}+4}}$, we get
$\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|-\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|g_{x}(s)|\cdot\left\|(M-sI)^{-1}\right\|\,|\mathrm{d}s|$
$\leq$ $\frac{1}{2\pi}\lim_{\begin{subarray}{c}\epsilon\to 0\end{subarray}}\int_{-R}^{-\epsilon}|g_{x}(s_{1}+\delta i)|\cdot\left\|(M-(s_{1}+\delta i)I)^{-1}\right\|\mathrm{d}s_{1}$
$-\frac{1}{2\pi}\lim_{\begin{subarray}{c}\epsilon\to 0\end{subarray}}\int^{-R}_{-\epsilon}|g_{x}(s_{1}-\delta i)|\cdot\left\|(M-(s_{1}-\delta i)I)^{-1}\right\|\mathrm{d}s_{1}$
$=$ $\frac{1}{\pi}\int_{0}^{R}|g_{x}(-s_{1})|\cdot\left\|(M+s_{1}I)^{-1}\right\|\mathrm{d}s_{1}$
$\leq$ $\frac{R^{2}}{\pi\sigma_{\min}(M)(\sqrt{R^{2}+x^{2}})}$
$=$ $\frac{4\sigma_{\max}(M)\kappa(M)}{\pi(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}.$
Finally, combining all estimates, we obtain the upper bound:
(14) $\left\|h_{-x}(M)\right\|\leq\frac{1+\pi}{\pi}\frac{4\sigma_{\max}(M)\kappa(M)}{(\sqrt{4(\sigma_{\max}(M))^{2}+x^{2}})}\leq\frac{4(1+\pi)\kappa(M)}{\pi}\frac{\sigma_{\max}(M)}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})}.$
Step 3: Apply Beta and Gamma Functions. Substituting (14) into (11) gives
@@
$\leq\left(\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\cdot\left\|g_{x}(M)\right\|_{\mathcal{S}}^{k}\cdot\left\|h_{-x}\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$\leq\frac{4(1+\pi)\kappa(M)}{\pi^{2}}\left(\int_{0}^{\infty}x^{1/2}\cdot\left\|g_{x}\right\|_{\mathcal{S}}^{k}\cdot\frac{\sigma_{\max}(M)}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$\leq\frac{4(1+\pi)\kappa(M)}{\pi^{2}}\left(\int_{0}^{\infty}x^{1/2}\cdot\frac{(\sigma_{\max}(M))^{k+1}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k+1}}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Define $\mathrm{I}:=\int_{0}^{\infty}x^{1/2}\cdot\frac{(\sigma_{\max}(M))^{k+1}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k+1}}\,\mathrm{d}x$. Let $x=\sigma_{\max}(M)y$, then we obtain
$\rm{I}$ $=(\sigma_{\max}(M))^{3/2}\int^{\infty}_{0}\frac{y^{1/2}}{(\sqrt{1+y^{2}})^{k+1}}\mathrm{~{}d}y$
$=(\sigma_{\max}(M))^{3/2}\int^{\frac{\pi}{2}}_{0}\frac{(\tan(\theta))^{1/2}(\sec(\theta))^{2}}{(\sec(\theta))^{k+1}}\mathrm{~{}d}\theta$
$=(\sigma_{\max}(M))^{3/2}\int^{\frac{\pi}{2}}_{0}(\sin(\theta))^{1/2}(\cos(\theta))^{k-3/2}\mathrm{~{}d}\theta$
$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\mathrm{B}\left(\frac{3}{4},\frac{k}{2}-\frac{1}{4}\right)$
$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\frac{\Gamma(3/4)\Gamma(k/2-1/4)}{\Gamma(k/2+1/2)}.$
Here, $\text{B}(\cdot,\cdot)$ and $\Gamma(\cdot)$ are the standard Beta and Gamma functions, respectively, with $\mathrm{B}(x,y)=\frac{\Gamma(x)\Gamma(y)}{\Gamma(x+y)}.$ Furthermore, noting that in [36] for $w>0$ and $l\in(0,1)$,
$\left(\frac{w}{w+l}\right)^{1-l}\leq\frac{\Gamma(w+l)}{w^{l}\Gamma(w)}\leq 1,$
we arrive at
$\frac{\Gamma(k/2-1/4)}{\Gamma(k/2+1/2)}$ $\leq\frac{(k/2+1/2)^{1/4}}{(k/2-1/4)^{1/4}(k/2-1/4)^{3/4}}$
$\leq\frac{2^{5/4}}{(k-1/2)^{3/4}}$
$\leq\frac{\sqrt{2}\pi^{2}}{(1+\pi)\Gamma(3/4)(k-1/2)^{3/4}}.$
Then, it holds that
$\mathrm{I}\leq\frac{\pi^{2}(\sigma_{\max}(M))^{3/2}}{\sqrt{2}(1+\pi)(k-1/2)^{3/4}}.$
Collecting all estimates leads to
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\frac{2\sqrt{2}\kappa(M)(\sigma_{\max}(M))^{3/2}}{(k-1/2)^{3/4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
Thus, the relative error of the Arnoldi approximation for matrix square roots is bounded by:
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}$ $\leq\frac{2\sqrt{2}\kappa(M)(\sigma_{\max}(M))^{3/2}}{(k-1/2)^{3/4}}\frac{\left\|\bm{\xi}_{0}^{k}\right\|}{\|M^{1/2}\bm{b}\|}$
(15) $\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}.$
Step 4: Bound $\|\bm{r}^{k}_{0}\|$. From (6), it holds that
$\bm{r}^{k}_{0}$ $=\bm{b}-MQ_{k}H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}$
$=\bm{b}-(Q_{k}H_{k}+h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T})H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}$
$=-h_{k+1,k}\bm{q}_{k+1}\bm{e}_{k}^{T}H_{k}^{-1}\|\bm{b}\|\bm{e}_{1}.$
Taking the 2-norm on both sides gives
(16) $\left\|\bm{r}^{k}_{0}\right\|\leq h_{k+1,k}\left\|H_{k}^{-1}\right\|\left\|\bm{b}\right\|\leq\frac{h_{k+1,k}}{\sigma_{\min}\left(H_{k}\right)}\left\|\bm{b}\right\|.$
Multiplying $\bm{e}_{k}$ on both sides of (6) and then taking the 2-norm, we have
(17) $h_{k+1,k}=\|M\bm{q}_{k}-Q_{k}H_{k}\bm{e}_{k}\|\leq\|M\bm{q}_{k}\|+\|H_{k}\bm{e}_{k}\|\leq\sigma_{\max}(M)+\sigma_{\max}(H_{k}),$
where the second inequality is achieved by utilizing the triangle inequality and the orthogonality of $Q_{k}$. Plugging the above estimate into (16) and utilizing the interlacing inequalities for singular values, we further bound $\left\|\bm{r}_{0}^{k}\right\|$ by
$\left\|\bm{r}^{k}_{0}\right\|\leq\frac{\sigma_{\max}(M)+\sigma_{\max}(H_{k})}{\sigma_{\min}\left(H_{k}\right)}\left\|\bm{b}\right\|\leq\frac{2\sigma_{\max}\left(M\right)}{\sigma_{\min}\left(M\right)}\left\|\bm{b}\right\|=2\kappa(M)\|\bm{b}\|.$
Consequently, substituting the above estimate of $\bm{r}_{0}^{k}$ into (3.9), we obtain
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{2\sqrt{2}(\kappa(M))^{5/2}}{(k-1/2)^{3/4}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{4\sqrt{2}(\kappa(M))^{7/2}}{(k-1/2)^{3/4}}.$
@@
If the matrix $M$ is Hermitian, then using the results from [6] and [1], we obtain
$\frac{\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|}{\|M^{1/2}\bm{b}\|}\leq\frac{(\kappa(M))^{3/2}}{2k^{3/2}}\frac{\|\bm{r}^{k}_{0}\|}{\|\bm{b}\|}\leq\frac{(\kappa(M))^{5/2}}{k^{3/2}}.$
This sharper result stems from the fact that the Hermitian matrix is unitarily diagonalizable with real eigenvalues, which allows for a more refined spectral analysis of $\left|\operatorname{det}\left(h_{z}\left(H_{k}\right)\right)\right|$ and $\left\|h_{z}(M)\right\|$ in 3.9.
@@
Note that the residual $\bm{r}_{0}^{k}$ is precisely the residual from the FOM applied to the linear system $M\bm{x}=\bm{b}$. Since the error in Theorem 3.8 is bounded in terms of $\bm{r}_{0}^{k}$, which is naturally computed in the Arnoldi iteration, we thus have a practical stopping criterion for our algorithm.

### v3

@@ 3.2 Matrix square root
$+\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{+}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|$
@@
$-\left(\frac{1}{2\pi}\int_{\mathcal{C}_{\delta}^{-}}|z^{1/2}|\prod_{i=1}^{k}\left|\frac{\lambda_{i}\left(H_{k}\right)}{\lambda_{i}(H_{k})-z}\right|\left\|(M-zI)^{-1}M\right\|\,|\mathrm{d}z|\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
As the outer radius $R\to\infty$, the integral over the large circular arc vanishes for $k\geq 1$ since $|z^{1/2}|=\mathcal{O}\left(R^{1/2}\right)$, $\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})-z}\right|=\mathcal{O}(R^{-1})$, and $\|M(M-zI)^{-1}\|=\mathcal{O}(R^{-1})$. Similarly, as the inner radius $\epsilon\to 0$, the integral goes to zero. Thus, as $\epsilon\to 0$ and $R\to\infty$, the contributions to the integral come from $\mathcal{C}^{+}_{\delta}$ and $\mathcal{C}^{-}_{\delta}$. Also, as $\epsilon\to 0$, $\delta$ approaches 0. Then, we have
@@
$\leq$ $\lim_{\begin{subarray}{c}\epsilon\to 0\\ R\to\infty\end{subarray}}\left(\frac{1}{2\pi}\int_{-R}^{-\epsilon}\left|(x+\delta i)^{1/2}\right|\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})-(x+\delta_{i})}\right|\left\|\left(M-(x+\delta iI)\right)^{-1}M\right\|\mathrm{~d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$-\left(\frac{1}{2\pi}\int^{-R}_{-\epsilon}\left|(x-\delta i)^{1/2}\right|\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})-(x+\delta_{i})}\right|\left\|\left(M-(x-\delta iI)\right)^{-1}M\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
$=$ $\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\left\|(M+xI)^{-1}M\right\|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|.$
Next, we prove that $\left\|(M+xI)^{-1}M\right\|<1$. This is equivalent to showing that for any non-zero vector $v\in\mathbb{C}^{n}$ and $x>0$,
$\left\|(M+xI)^{-1}M\bm{v}\right\|<\left\|\bm{v}\right\|.$
To prove this, let $\bm{u}=(M+xI)^{-1}M\bm{v}$, and note that $M(\bm{u}-\bm{v})=-x\bm{u}.$ Taking the inner product with $\bm{u}-\bm{v}$, we have
$x\operatorname{Re}\left(\bm{u},\bm{u}-\bm{v}\right)=\operatorname{Re}\left(M(\bm{u}-\bm{v}),\bm{u}-\bm{v}\right)>0.$
Thus, we have
$0>\operatorname{Re}\left(\bm{u},\bm{u}-\bm{v}\right)=\left\|\bm{u}\right\|^{2}-\operatorname{Re}\left(\bm{u},\bm{v}\right)\geq\left\|\bm{u}\right\|^{2}-\|\left(\bm{u},\bm{v}\right)\|\geq\left\|\bm{u}\right\|^{2}-\left\|\bm{u}\right\|\left\|\bm{v}\right\|$
which implies $\left\|\bm{u}\right\|<\left\|\bm{v}\right\|$ and, therefore, $\left\|(M+xI)^{-1}M\right\|<1$. Thus,
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|=\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|,$
which completes the proof.
@@
In the Arnoldi process, $H_{k}$ is generated naturally, and the FOM error $\left\|\bm{\xi}_{0}^{k}\right\|$ can be obtained with essentially one additional step. Therefore, the error bound in Theorem 3.8 provides an a posteriori error estimate that can be computed efficiently in practice, especially for small $k$. This motivates the following practical stopping criterion:
$\frac{1}{\pi}\left\|\bm{\xi}_{0}^{k}\right\|\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\leq\textit{tol}.$
where tol is a given tolerance.
Next, building on the a posteriori error estimate in Theorem 3.8, we derive an a priori error estimate, which is summarized in the following theorem.
Theorem 3.11.
Given a positive-definite matrix $M\in\mathbb{C}^{n\times n}$ and a vector $\bm{b}\in\mathbb{C}^{n}$, for $k\geq 2$, the error of the Arnoldi process for approximating the matrix square root is bounded by
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\frac{\Gamma\!(3/4)}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|,$
where $\Gamma\!(\cdot)$ is the standard Gamma function, and $\sigma_{\max}(M)$ denotes the largest singular value of $M$.
Proof 3.12.
From Theorem 3.8, we know that
$\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|\leq\frac{1}{\pi}\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x.$
Consider any $w\in\mathbb{C}$ with $\operatorname{Re}(w)>0$ and fix $x>0$ in $\mathbb{R}$, a straightforward calculation then yields
$\left|\frac{w}{w+x}\right|\leq\frac{|w|}{\sqrt{|w|^{2}+x^{2}}}.$
The function on the right-hand side is increasing with respect to $|w|$. Denote $\sigma_{\min}(M)$ as the smallest eigenvalue of $M$. Applying the interlacing inequalities for singular values in Theorem 2 of [thompson1972principal] gives
$\max_{i}|\lambda_{i}(H_{k})|\leq\sigma_{\max}(H_{k})\leq\sigma_{\max}(M),\qquad\min_{i}|\lambda_{i}(H_{k})|\geq\sigma_{\min}(H_{k})\geq\sigma_{\min}(M).$
Letting $\mathcal{S}:=[\sigma_{\min}(M),\sigma_{\max}(M)]$, one arrives at
$\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x$ $\leq\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{|\lambda_{i}(H_{k})|}{\sqrt{|\lambda_{i}(H_{k})|^{2}+x^{2}}}\,\mathrm{d}x$
$\leq\int_{0}^{\infty}x^{1/2}\left(\sup\limits_{|w|\in\mathcal{S}}\frac{|w|}{\sqrt{|w|^{2}+x^{2}}}\right)^{k}\,\mathrm{d}x$
$\leq\int_{0}^{\infty}x^{1/2}\frac{(\sigma_{\max}(M))^{k}}{(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}})^{k}}\,\mathrm{d}x.$
Letting $x=\sigma_{\max}(M)\tan(\theta)$, for $k\geq 2$, we obtain
$\int_{0}^{\infty}x^{1/2}\frac{(\sigma_{\max}(M))^{k}}{\bigl(\sqrt{(\sigma_{\max}(M))^{2}+x^{2}}\bigr)^{k}}\,\mathrm{d}x$ $=(\sigma_{\max}(M))^{3/2}\int_{0}^{\frac{\pi}{2}}\frac{\tan^{1/2}(\theta)\,\sec^{2}(\theta)}{\sec^{k}(\theta)}\,\mathrm{d}\theta$
$=(\sigma_{\max}(M))^{3/2}\int_{0}^{\frac{\pi}{2}}\sin^{1/2}(\theta)\,\cos^{k-5/2}(\theta)\,\mathrm{d}\theta$
$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\,\mathrm{B}\!\left(\frac{3}{4},\,\frac{2k-3}{4}\right)$
$=\frac{(\sigma_{\max}(M))^{3/2}}{2}\,\frac{\Gamma\!\left(3/4\right)\Gamma\!\left((2k-3)/4\right)}{\Gamma\!\left(k/2\right)},$
where $\mathrm{B}(\cdot,\cdot)$ is the Beta function, satisfying $\mathrm{B}(x,y)=\frac{\Gamma(x)\Gamma(y)}{\Gamma(x+y)}$. Furthermore, following [wendel1948note], for $w>0$ and $l\in(0,1)$,
$\frac{w}{(w+l)^{1-l}}\leq\frac{\Gamma(w+l)}{\Gamma(w)}\leq w^{l},$
Then, with w=(2k3)/4w=(2k-3)/4 and $l=3/4$, we have,
$\frac{\Gamma\!\left((2k-3)/4\right)}{\Gamma\!\left(k/2\right)}\leq\frac{4}{2k-3}\left(\frac{k}{2}\right)^{\frac{1}{4}}=2^{\frac{7}{4}}\frac{k^{1/4}}{2k-3},$
Collecting all estimates yields the desired result.
@@
Summarizing the above results, we obtain the following estimates:
(9) $\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\bm{b})\|$ $\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\left|\frac{\lambda_{i}(H_{k})}{\lambda_{i}(H_{k})+x}\right|\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
(10) $\leq\frac{1}{\pi}\left(\int_{0}^{\infty}x^{1/2}\prod_{i=1}^{k}\frac{|\lambda_{i}(H_{k})|}{\sqrt{|\lambda_{i}(H_{k})|^{2}+x^{2}}}\,\mathrm{d}x\right)\left\|\bm{\xi}_{0}^{k}\right\|$
(11) $\leq\frac{\Gamma\!(3/4)}{2^{1/4}\pi}\frac{2k}{2k-3}(\sigma_{\max}(M))^{\frac{3}{2}}k^{-\frac{3}{4}}\left\|\bm{\xi}_{0}^{k}\right\|.$
The inequalities Eq. 9 and Eq. 10 provide a posteriori error bounds and remain relatively tight. The last inequality Eq. 11 gives an a priori bound; since it relies on $|\lambda_{i}(H_{k})|\leq\sigma_{\max}(M)$, it becomes loose when the eigenvalues of $M$ are widely distributed.